# Neural Network-based Universal Formulas for Control

### v1

@@ 5 Simulations
In this section we illustrate our approach in different simulation examples. We train a NN to approximate $k^{*}$ as detailed in Section 4. We focus on the case of $m=2$ and $N=2$. We use a feedforward NN with 4 layers (with input dimension $7$ and output dimension $2$), each with 64 neurons, and we use the Sigmoid Linear Unit (SiLU) activation function $\phi:\mathbb{R}\to\mathbb{R}$ given by $\phi(s)=\frac{s}{1+\exp\{-s\}}$ (Elfwing et al., 2018), which is a smooth approximation of the ReLU activation function. The NN is trained with the Adam optimizer with learning rate $3\times 10^{-4}$ for $2000$ epochs using a dataset with $25113$ data points. Each data point ${\mathbf{q}}$ consists of a uniformly randomly sampled point in $\left([-1,1]\times\mathcal{B}_{1}(\mathbf{0}_{2})\times[-1,1]\times\mathcal{B}_{1}(\mathbf{0}_{2})\times[0,1]\right)\cap\mathcal{P}\subset\mathbb{R}^{7}$ (to do so, we first uniformly sample a point from $[-1,1]\times\mathcal{B}_{1}(\mathbf{0}_{2})\times[-1,1]\times\mathcal{B}_{1}(\mathbf{0}_{2})\times[0,1]$ and then check whether the corresponding 2 inequalities are feasible with the QP solver of the cvxpy library in Python (Diamond and Boyd, 2016)), along with the corresponding minimizer of the function $\tilde{J}_{{\mathbf{q}}}$. We compute such minimizer up to a tolerance of $10^{-6}$ by numerically integrating the gradient flow $\dot{{\mathbf{k}}}=-\nabla\tilde{J}_{{\mathbf{q}}}({\mathbf{k}})$ using Python’s solve_ivp function in the SciPy library (Virtanen et al., 2020). We also train the same NN architecture with the HardNet method from (Min and Azizan, 2025), which includes a projection step at the output layer and guarantees that the NN predictions satisfy the two affine constraints (cf. Remark 4.4). All simulations were run in an Ubuntu PC with Intel Core i9-13900K 3 GHz 24-Core Processor.
@@
(Safe stabilization of single-integrator system): Consider a single integrator on $\mathbb{R}^{2}$,
$\dot{x}=u_{x},\qquad\dot{y}=u_{y}.$
@@
Suppose that our goal is to design a controller that stabilizes the system to the origin and stays in the safe set $\mathcal{C}=\{(x,y)\in\mathbb{R}^{2}:h(x,y):=x^{2}+(y-2.5)^{2}-1\geq 0\}$. Note that $V(x,y)=\frac{1}{2}(x^{2}+y^{2})$ is a CLF and $h$ is a CBF of $\mathcal{C}$. Take the positive definite function $W$ in (2) to be $W(x,y)=0.1(x^{2}+y^{2})$ and the extended class $\mathcal{K}_{\infty}$ function $\alpha$ in (3) to be $\alpha(s)=s$. Now, using the inequalities (2), (3), we define $a_{1}(x,y)=0.1(x^{2}+y^{2})$, $b_{1}(x,y)=(x,y)$, $a_{2}(x,y)=-h(x,y)$, and $b_{2}(x,y)=(-2x,-2(y-2.5))$. By using (Mestres and Cortés, 2023, Lemma 5.2), we can show that the inequalities $a_{1}(x,y)+b_{1}(x,y)^{\top}{\mathbf{u}}<0$, $a_{2}(x,y)+b_{2}(x,y)^{\top}{\mathbf{u}}<0$ are simultaneously feasible for $\{(x,y)\in\mathbb{R}^{2}:x\neq 0\ \text{or}\ y\leq 1.3\}\subset\mathcal{F}$. Figure 1 compares the trajectories obtained by directly applying the NN-based controller in closed-loop and using it to warmstart an optimization scheme to compute $u^{*}$ online. We observe that using the NN-based controller directly fails to exactly stabilize the origin (trajectories converge to a neighborhood of it) and induces safe trajectories for the chosen initial conditions. Instead, the trajectories that use a controller obtained by numerically computing $u^{*}$ online with the NN controller as a warmstart asymptotically converge to the origin and are also safe. The same is true for the controllers obtained from the HardNet method, the CLF-CBF QP, or the computation of $u^{*}$ online with the CLF-CBF QP as a warmstart.
@@
Figure 1: (left) Trajectories of the closed-loop system obtained from the neural network based controller for Example 5.1. (right) Trajectories of the closed-loop system obtained from numerically finding the controller $u^{*}$ online and warmstarting the solver with the NN-based controller for Example 5.1. Initial conditions are denoted by red crosses, the origin is the black dot, and the green region denotes the unsafe set.
@@
Table 1 reports the execution times of the various considered controllers. The NN-based controller is the fastest, but does not possess safety and stability guarantees. The HardNet-based controller is one order of magnitude slower but has safety and stability guarantees by construction, although no optimality guarantees. The CLF-CBF QP controller and the controllers obtained by computing $u^{*}$ online are two orders of magnitude slower than the NN-based controller, but are safe, stable, and optimal by construction up to numerical errors. It is also worth noting that warmstarting the numerical solver to optimize $\tilde{J}_{{\mathbf{q}}}$ with the NN prediction also improves its computational time compared to warmstarting it with the CLF-CBF QP.
@@
Table 1: Average execution times ($\pm$ standard deviation) in milliseconds over the trajectories in Figure 1 for different controller implementations. The first column refers to the controller obtained directly as the prediction of the NN. The second column refers to the controller obtained as the prediction of the NN when trained with the HardNet method from (Min and Azizan, 2025). The third column refers to the controller obtained by solving the CLF-CBF QP from (Ames et al., 2017) (using the cvxpy library in Python (Diamond and Boyd, 2016)). The fourth (resp. fifth) column refers to the controller obtained by finding the minimum of the function $\tilde{J}_{{\mathbf{q}}}$ using a numerical solver (the solve_ivp function in Python’s SciPy library) and warmstarting the solver with the CLF-CBF QP (resp. the NN prediction).
@@
with ${\mathbf{u}}=(v,\omega)$ being the control input. Suppose that our goal is to design a controller that stabilizes the system to the origin and stays in the safe set $\mathcal{C}=\{(x,y,\theta)\in\mathbb{R}^{3}:h(x,y,\theta)=-y+(2x+1)^{2}+1\geq 0\}$ (this example is taken from (Ong and Cortés, 2019, Section VII). Take the positive definite function $W(x,y)=0.1(x^{2}+y^{2}+\theta^{2})$ and $\alpha(s)=s$. Now, using the inequalities (2), (3), we can define $b_{1}(x,y,\theta)=(x\cos(\theta)+y\sin(\theta),\theta)$ $a_{1}(x,y,\theta)=-y^{2}+0.1(x^{2}+y^{2}+\theta^{2})$, $b_{2}(x,y,\theta)=(-4(2x+1)\cos(\theta)+\sin(\theta),0)$, $a_{2}(x,y,\theta)=-y-2h(x,y)$. As shown in (Ong and Cortés, 2019, Section VII), the inequalities $a_{1}(x,y,\theta)+b_{1}(x,y,\theta)^{\top}{\mathbf{u}}<0$, $a_{2}(x,y,\theta)+b_{2}(x,y,\theta)^{\top}{\mathbf{u}}<0$ are simultaneously feasible in all of $\mathcal{C}\backslash\{\mathbf{0}_{2}\}$, so $\mathcal{C}\backslash\{\mathbf{0}_{2}\}\subset\mathcal{F}$. Since the dimension of the input is $2$ and we have $2$ constraints, we can use the same NN as in Example 5.1, even though the state dimension is different. Figure 2 compares the trajectories obtained by directly applying the NN-based controller in closed-loop and using it to warmstart an optimization scheme to compute $u^{*}$ online. In this case, both controllers induce safe trajectories that asymptotically converge to the origin. Table 1 reports the execution times of various controllers for this example as well.
Figure 2: (left) Projection in the $(x,y)$ plane of trajectories of the closed-loop system obtained from the neural network based controller for Example 5.2. (right) Projection in the $(x,y)$ plane of trajectories of the closed-loop system obtained from numerically finding the controller $u^{*}$ online and warmstarting the solver with the NN-based controller for Example 5.2. Initial conditions are denoted by red crosses (and all have an initial orientation $\theta_{0}=\pi+0.1$), the origin is the black dot, and the green region denotes the unsafe set. Black arrows indicate the orientation of the unicycle (i.e., the $\theta$ variable) at that point of the trajectory. Observe that the velocity $v$ could be negative, so that at points near the right of the target, the vehicle is "backing up".
@@ 6 Conclusions, Limitations, and Future Work
We have studied the problem of designing a controller that satisfies an arbitrary number of affine inequalities at every point in the state space. This is an important problem in constructing guardrails for autonomous systems, since safety, stability, and input constraints can be achieved by controllers satisfying such inequalities. We have provided a novel universal formula for controllers satisfying such constraints. The controller is given at every state as the minimizer of a strictly convex function. To avoid the computation of such minimizer in real time, we have introduced a method based on NN to approximate it. Remarkably, this NN is universal in the sense that it can be used for any control task with input dimension and number of constraints less than some fixed value, and can be trained with data from just a small subset of the state space. We have shown the performance of the controller and its NN approximation in various simulation examples. A limitation of the presented approach is that it has not been tested in systems with high-dimensional states and inputs, or a high number of constraints. Additionally, the presented controller is only defined in the region where the constraints are simultaneously feasible, and is limited to affine constraints. Future work will focus on addressing these shortcomings.

### v2

@@ 5 Simulations
In this section we illustrate our approach in different simulation examples. We train a NN to approximate $k^{*}$ as detailed in Section 4. We focus on the case of $m=2$ and $N=2$. We use a feedforward NN with 4 layers (with input dimension $7$ and output dimension $2$), each with 64 neurons, and we use the Sigmoid Linear Unit (SiLU) activation function $\phi:\mathbb{R}\to\mathbb{R}$ given by $\phi(s)=\frac{s}{1+\exp\{-s\}}$ [37], which is a smooth approximation of the ReLU activation function. The NN is trained with the Adam optimizer with learning rate $3\times 10^{-4}$ for $2000$ epochs using a dataset with $25113$ data points. Each data point $\mathbf{q}$} consists of a uniformly randomly sampled point in $\left([-1,1]\times\mathcal{B}_{1}(\mathbf{0}_{2})\times[-1,1]\times\mathcal{B}_{1}(\mathbf{0}_{2})\times[0,1]\right)\cap\mathcal{P}\subset\mathbb{R}^{7}$ (to do so, we first uniformly sample a point from $1,1]\times\mathcal{B}_{1}(\mathbf{0}_{2})\times[-1,1]\times\mathcal{B}_{1}(\mathbf{0}_{2})\times[0,1]$ and then check whether the corresponding 2 inequalities are feasible with the QP solver of the cvxpy library in Python [38]), along with the corresponding minimizer of the function $\tilde{J}_{{\mathbf{q}}}$. We compute such minimizer up to a tolerance of $10^{-6}$ by numerically integrating the gradient flow $\dot{{\mathbf{k}}}=-\nabla\tilde{J}_{{\mathbf{q}}}({\mathbf{k}})$ using Python’s solve_ivp function in the SciPy library [39]. We also train the same NN architecture with the HardNet method from [36], which includes a projection step at the output layer and guarantees that the NN predictions satisfy the two affine constraints (cf. Remark 4.4). All simulations were run in an Ubuntu PC with Intel Core i9-13900K 3 GHz 24-Core Processor.
@@
(Safe stabilization of single-integrator system): Consider a single integrator on $\mathbb{R}^{2}$, i.e., $\dot{x}=u_{x},\dot{y}=u_{y}$. Suppose that our goal is to design a controller that stabilizes the system to the origin and stays in the safe set $\mathcal{C}=\{(x,y)\in\mathbb{R}^{2}:h_{1}(x,y)=x^{2}+(y-2.5)^{2}-1\geq 0,\ h_{2}(x,y)=(x+2)^{2}+(y+2)^{2}-1\geq 0,\ h_{3}(x,y)=(x-2)^{2}+(y+2)^{2}-1\geq 0\}$. Note that $V(x,y)=\frac{1}{2}(x^{2}+y^{2})$ is a CLF and $h(x,y):=h_{1}(x,y)h_{2}(x,y)h_{3}(x,y)$ is a CBF of $\mathcal{C}$. Take the positive definite function $W$ in (2) to be $W(x,y)=0.1(x^{2}+y^{2})$ and the extended class $\mathcal{K}_{\infty}$ function $\alpha$ in (3) to be $\alpha(s)=s$. Now, using the inequalities (2), (3), we define $a_{1}(x,y)=0.1(x^{2}+y^{2})$, $b_{1}(x,y)=(x,y)$, $a_{2}(x,y)=-h(x,y)$, and $b_{2}(x,y)=-\nabla h(x,y)$. By using [13, Lemma 5.2], we can show that the inequalities $a_{1}(x,y)+b_{1}(x,y)^{\top}{\mathbf{u}}<0$, $a_{2}(x,y)+b_{2}(x,y)^{\top}{\mathbf{u}}<0$ are simultaneously feasible at all points in $\mathcal{C}$ not satisfying $\frac{\partial h}{\partial x}(x,y)y=\frac{\partial h}{\partial y}(x,y)x$. More explicitely, the inequalities will be satisfied if there is a pair $(u_{x},u_{y})$ so that the entries of
$\begin{pmatrix}a_{1}(x,y)\\ a_{2}(x,y)\end{pmatrix}\;+\;\begin{pmatrix}x&y\\ -\frac{\partial h}{\partial x}(x,y)&-\frac{\partial h}{\partial y}(x,y)\end{pmatrix}\begin{pmatrix}u_{x}\\ u_{y}\end{pmatrix}$
@@
are both negative, which can be achieved unless the determinant of the matrix is zero. Many points in the set where the determinant is zero are also in $\mathcal{F}$, but this is sufficient to show that $\mathcal{F}$ consists of all points in $\mathcal{C}$ except for a set of measure zero. Figure 1 compares the trajectories obtained by directly applying the NN-based controller in closed-loop and using it to warmstart an optimization scheme to compute $u^{*}$ online. We observe that the NN-based controller induces trajectories that are safe and converge to a small neighborhood around the origin, for the chosen initial conditions. Instead, the trajectories that use a controller obtained by numerically computing $u^{*}$ online with the NN controller as a warmstart asymptotically converge to the origin (instead of a neighborhood of it) and are also safe. The same is true for the controllers obtained from the HardNet method, the CLF-CBF QP, or the computation of $u^{*}$ online with the CLF-CBF QP as a warmstart.
@@
Figure 1: (left) Trajectories of the closed-loop system obtained from the neural network based controller for Example 5.1. (right) Trajectories of the closed-loop system obtained from numerically finding the controller $u^{*}$ online, warmstarting the solver with the NN-based controller for Example 5.1. Initial conditions are denoted by red crosses, the origin is the black dot, and the green region denotes the unsafe set.
@@
Table 1, top row reports the execution times of the various considered controllers. The NN-based controller is the fastest, but does not possess safety and stability guarantees. The HardNet-based controller is one order of magnitude slower but has safety and stability guarantees by construction, although no optimality guarantees. The CLF-CBF QP controller and the controllers obtained by computing $u^{*}$ online are two orders of magnitude slower than the NN-based controller, but are safe, stable, and optimal by construction up to numerical errors. It is also worth noting that warmstarting the numerical solver to optimize $\tilde{J}_{{\mathbf{q}}}$ with the NN prediction also improves its computational time compared to warmstarting it with the CLF-CBF QP.  $\bullet$
@@
Table 1: Average execution times ($\pm$ standard deviation) in milliseconds over the trajectories in Figure 1 (top row) and Figure 2 (bottom row) for different controller implementations. The first column refers to the controller obtained directly as the prediction of the NN. The second column refers to the controller obtained as the prediction of the NN when trained with the HardNet method from [36]. The third column refers to the controller obtained by solving the CLF-CBF QP from [6] (using the cvxpy library in Python [38]). The fourth (resp. fifth) column refers to the controller obtained by finding the minimum of the function $\tilde{J}_{{\mathbf{q}}}$ using a numerical solver (the solve_ivp function in Python’s SciPy library) and warmstarting the solver with the CLF-CBF QP (resp. the NN prediction).
@@
[…]
Figure 2: (left) Projection in the $(x,y)$ plane of trajectories of the closed-loop system obtained from the neural network based controller for Example 5.2. (right) Projection in the $(x,y)$ plane of trajectories of the closed-loop system obtained from numerically finding the controller $u^{*}$ online and warmstarting the solver with the NN-based controller for Example 5.2. Initial conditions are denoted by red crosses (and all have an initial orientation $\theta_{0}=\pi+0.1$), the origin is the black dot, and the green region denotes the unsafe set. Black arrows indicate the orientation of the unicycle (i.e., the $\theta$ variable) at that point of the trajectory. Observe that the velocity $v$ could be negative, so that at points near the right of the target, the vehicle is "backing up". We note also that due to the presence of a drift term in the $y$ dynamics, the orientation of the unicycle might not be tangent to the trajectory.
@@ 6 Conclusions, Limitations, and Future Work
We have studied the problem of designing a controller that satisfies an arbitrary number of affine inequalities at every point in the state space, which arise when enforcing safety, stability, and input constraints. We have provided a novel universal formula for controllers satisfying such affine inequalities. The control input is given at every state as the minimizer of a strictly convex function. To avoid the computation of such minimizer in real time, we have introduced a method based on NN to approximate it. Remarkably, this NN is universal in the sense that it can be used for any control task with input dimension and number of constraints less than some fixed value, and can be trained with data from just a small subset of the state space. We have shown the performance of the controller and its NN approximation in various simulation examples. Future work will test our proposed approach to systems with high-dimensional states and inputs, or a high number of constraints. We also plan to extend the proposed methodology to non-affine constraints.